Graph without a calculator:
1) 3x - 2y = 8
2) 4x + 5y = 3
I know how to plot equation one using y = mx + b and I know the intersection point is (2, -1). However, equation two is giving me a headache and I have no idea how to plot (-4x + 3)/5, which has the same denominator.

Found 2 solutions by Alan3354, edjones:
Answer by Alan3354(69443)

Graph without a calculator:
1) 3x - 2y = 8
2) 4x + 5y = 3
I know how to plot equation one using y = mx + b and I know the intersection point is (2, -1). However, equation two is giving me a headache and I have no idea how to plot (-4x + 3)/5, which has the same denominator.
--------------
4x + 5y = 3
y = (-4/5)x + 3/5
Can you do that?

Answer by edjones(8007)

These are linear equations, straight lines.
Use graph paper or the equivalent.
y=(-4x + 3)/5
When x=0 y=3/5 Make a dot at this point on the graph paper.
When x=10 y= -37/5 = -7 2/5 Make another dot at this point.
Connect the 2 dots.
I know the dots are not on any of the lines but between them and you can't be exact. It should be close enough.
.
Ed